An organ pipe contains oxygen gas at 47C resonates at a frequency of 300Hz,if the pipe is now filled with nitrogen gas then for the organ pipe to resonate at 300Hz what should be the temperature? 
  1. 7
  2. 41.1
  3. 280
  4. 92.7

Grade:12

1 Answers

noogler
8 years ago
n1=n2
n is proportional to c here we take rms velcity
V1=V2
Vrms=under root 3RT/M
T1/M1=T2/M2
320/32=T/28
T=280K=7OC